Find mean and margin of error if    28.4 < µ < 38.8Find mean and margin of error if    28.4 < µ < 38.8
Answer #1

solution:

upper limit=38.8

lower limit=28.4

mean=upper limit+lower limit/2

=38.8+28.4/2

=67.2/2

=33.6

margin of error=upper limit-lower limit/2

=38.8-28.4/2

=10.4/2

=5.2


answered by: ANURANJAN SARSAM